OpenStudy (anonymous):

standard form : parallel to 3x-6y =5 , containing (-2,4)

OpenStudy (anonymous):

\begin{array}c&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&\color{}{\bullet}&&&&\color{}{\bullet}&&&&&&&&&&&&&&&\\&\color{}{\bullet}&&&&\color{}{\bullet}&&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&\\&\color{}{\bullet}&&&&\color{}{\bullet}&&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&\\&\color{}{\bullet}&\color{}{\bullet}&\color{}{\bullet}&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&&&&\\&\color{}{\bullet}&&&&\color{}{\bullet}&&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&\\&\color{}{\bullet}&&&&\color{}{\bullet}&&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&\\&\color{}{\bullet}&&&&\color{}{\bullet}&&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&\\&&&&&&&\color{}{\bullet}&\color{}{\bullet}&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\&&&&&&&&&&&&&&&&&&&&\\\end{array}

OpenStudy (anonymous):

3x-6y=5 -6y=-3x+5 y=1/2*x+5/6 parallel means same slope, so y=1/2x+b plug in known values 4=1/2*-2+b 4=-1+b 5=b final, y=1/2x+5

OpenStudy (anonymous):

or, in the original form y=1/2x+5 2y=x+10 -x+2y=10

OpenStudy (aroub):

parallel =same slope -6y=-3x+5 y=1/2x-5/6 y=mx+b , m= 1/2 , (-2,4) 4=1/2(-2)+b 4+1=b 5=b y=1/2x+5

OpenStudy (anonymous):

thanks y'all(: